# Unveiling Super Experts in Mixture-of-Experts Large Language Models

### v1

@@ 3.2 Localization of Super Experts
Okay, so the problem is asking: How many of the first
one hundred positive integers are divisible by 3, 4,
and 5? Hmm, let me think. First, I need to understand
Prune Super Experts
Okay, let’s see here. So the question is how many of
the first one hundred positive integers are divisible
by 3,4,3,4,3 , 4 , and 555 right? Okay, so first, that’s probably
going to, like, but, that’s, the way, it’s, the way, it’s,
the way, it’s, the way, it’s, the way, it’s, the way, it’s,
… … Repeating … …
the way, it’s, the way, it’s, the way, it’s, the way, it’s
Random Pruning
Okay, so the problem is asking: How many of the first
one hundred positive integers are divisible by 3, 4,
and 5? Hmm, let me think. First, I need to understand
Table 6: Responses of DeepSeek-R1 in the MATH-500 (Lightman et al. 2023) benchmark.
(i) SEs are consistently present across the investigated models, accounting for less than 0.5% of all experts.
(ii) After post-training processes, the distribution of SEs remains unchanged compared to the base model.
(iii) Some experts in the final layers also exhibit extreme activation outliers. However, since they do not contribute to the formation of massive activations, they do not hold the same level of significance as SEs. Additional experimental results are available in Appendix A.
Distribution of SEs Across Different Data Domains
We also examine their patterns across various input data domains. Specifically, in addition to the C4 dataset (Raffel et al. 2020), we analyze SE distributions across several other datasets, including WikiText-2 (Merity et al. 2016), C-Eval (Huang et al. 2023), GSM8K (Cobbe et al. 2021), and HumanEval (Chen et al. 2021). As shown in the Appendix B, the distribution of SEs remains highly stable, regardless of variations in the input data domain.
4 The Importance of Super Experts
In this section, we evaluate the significance of SEs by quantifying the performance degradation across various capabilities following their pruning. We use the original model and the average results from random pruning of an equivalent number of experts as baselines. To more effectively evaluate the importance of SEs, we utilize distinct benchmark types for non-reasoning and reasoning models. For non-reasoning models, we conduct a series of few-shot evaluations, whereas for reasoning models, we employ more demanding benchmarks, encompassing tasks such as mathematical reasoning and coding.
4.1 Impact on Non-Reasoning Models
For non-reasoning models, we select three models: the non-thinking mode of Qwen3-30B-A3B, DeepSeek-V2-Lite and Mixtral-8x7B-v0.1. We utilize the datasets listed below and conduct evaluations using lm-eval (Gao et al. 2024), including ARC-challenge (ARC-c), ARC-easy (ARC-e) (Clark et al. 2018), BoolQ (Clark et al. 2019a), GSM8K (Cobbe et al. 2021), HellaSwag (Zellers et al. 2019), MMLU (Hendrycks et al. 2021), OpenBookQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), and WinoGrande (Keisuke et al. 2019). As shown in Table 3, pruning only a few super-experts leads to significant degradation across all tasks, with average accuracy dropping by 21.68% to 27.21%. In particular, for GSM8K, the degradation ranges from 52.71% to 74.15%. In contrast, random pruning has a negligible impact, underscoring the crucial role of SEs.
4.2 Impact on Reasoning Models
For evaluating the importance of SEs in reasoning models, we select DeepSeek-R1 and the thinking mode of Qwen3-30B-A3B. We select benchmarks more suitable for testing reasoning models and conduct evaluations based on the EvalScope (Team 2024a), a comprehensive model evaluation and performance benchmarking framework meticulously developed by the ModelScope community (Team 2023). The generation configurations used in the evaluation align with the corresponding technical reports of the models. These benchmarks are:
  • General Tasks: We use GPQA-Diamond under a 5-shot setting. GPQA (Rein et al. 2024) is a challenging dataset of multiple-choice questions authored by domain-specific multidisciplinary experts.
  • Math & Text Reasoning: To evaluate mathematical and logical reasoning skills, we use high-level math benchmarks, including MATH-500 (Lightman et al. 2023), AIME’24 (AIME 2024), and AIME’25 (AIME 2025).
  • Agent & Coding: To test the model’s proficiency in coding and agent-based tasks, we use LiveCodeBench (Jain et al. 2024) and HumanEval (Chen et al. 2021).
The experimental results, presented in Tables 4 and 5, show that pruning the SEs causes a significant performance degradation, while random pruning has almost no impact. The Pass@1 scores for most tasks drop to zero, highlighting the critical role of SEs. During the review of model responses on the Math-500 benchmark, we made a striking observation: after pruning the SEs, the model consistently generated repetitive responses in nearly every test, continuing until it reached the maximum output length, as shown in Table 6. This behavior suggests that the model loses its ability to reason and solve problems entirely after SE pruning. More results are in Appendix C.
(a) Layer 15 Head 10.
(b) Layer 25 Head 10.
(c) Layer 35 Head 10.
(d) Layer 15 Head 10.
(e) Layer 25 Head 10.
(f) Layer 35 Head 10.
Figure 6: Attention scores of Qwen3-30B-A3B. Figures (a), (b), and (c) depict the attention score maps of the original model, where the first token clearly functions as an attention sink, consistently attracting the majority of attention. Figures (d), (e), and (f) illustrate the attention scores following SE pruning, where the attention sink completely disappears.
Figure 7: Attention sink decay rate across layers.
5 Understanding the Impact of Super Experts Compression
In this section, we focus on deepening our understanding of the underlying impact of SE compression and further quantitatively assessing its effects.
5.1 Super Experts Induce Attention Sinks
Previous studies in dense LLMs (Sun et al. 2024; Gu et al. 2024; An et al. 2025) suggest that massive activations tend to attract disproportionately high attention scores in multi-head self-attention (MHSA) layers, leading to attention sinks for initial tokens. StreamLLM (Xiao et al. 2023b) conducted an initial investigation revealing the presence of attention sinks in LLMs. Subsequent studies (Gu et al. 2024; Guo et al. 2024) have demonstrated that although attention sinks often emerge at semantically insignificant tokens, the mechanism itself is critical for model performance. In efficient LLM techniques such as sparse attention and KV cache compression, maintaining attention sinks is essential for preventing undesirable distributional shifts of attention scores (Xiao et al. 2023b; Su et al. 2025). Building on this insight and our previous findings, we hypothesize that pruning SEs not only eliminates massive activations but also disrupts the underlying attention sink mechanism.
5.2 Impact of Super Experts Compression
To validate this insight and quantitatively assess the impact of SE pruning on attention sinks, we introduce (i) Attention Sink Decay Rate, denoted as $D_{\text{sink}}$. It is defined as the average decay rate of attention sinks for each head:
where $H$ is the total number of heads, $p_{i}^{t}$ represents the attention score between the Query token $t$ and the Key token $i$ before SE pruning, ${p_{i}^{t}}^{\prime}$ denotes the attention score after SE pruning, and $S$ refers to the set of sink tokens.
As shown in Figure 7, after pruning SEs, the attention sink decay rate remains consistently high, at approximately or even exceeding 90%, demonstrating a substantial disruptive effect on attention sinks. Figure 6 visualizes the attention scores for several heads before and after pruning SEs, highlighting the complete disappearance of attention sinks following SE pruning. Notably, attention sinks introduce implicit attention biases (Sun et al. 2024; An et al. 2025) that persist across all subsequent tokens and may encode global or other critical information (Darcet et al. 2023). Consequently, the impact of SE pruning on attention computation remains both continuous and significant.
6 Related Work
6.1 Expert-Level Compression
Beyond traditional LLM-focused compression methods, recent advancements have increasingly focused on expert-level compression strategies, capitalizing on the distinct structural features of MoE LLMs and the natural variations in expert importance. These approaches—including expert pruning, skipping, merging, and expert-level mixed-precision quantization—are being actively explored for their potential to enhance MoE model efficiency. M-SMoE (Li et al. 2023) performs expert merging by using activation frequencies to consolidate less significant experts, while also applying low-rank techniques to the merged experts to achieve further compression. NAEE (Lu et al. 2024) introduces plug-and-play pruning and skipping methods that leverage reconstruction loss to selectively compress less critical experts. MC (Huang et al. 2025) harnesses the significance of both experts and tokens to perform mixed-precision quantization and dynamic expert pruning, achieving extreme compression. MC-Suite (Jaiswal et al. 2025) reviews various empirical criteria for identifying critical experts, considering four dimensions: weight, expert behavior, intermediate activations, and gradient behavior. While these methods examine expert importance from various perspectives to optimize expert compression, they lack a deeper exploration and understanding of expert significance, overlooking the mechanistic importance of specific experts. Our work presents the first discovery and investigation of Super Experts, offering an in-depth analysis of their mechanistic role in inference, thereby filling a critical gap in current research.
7 Conclusion
In this work, we present the first identification and systematic study of a distinct subset of experts, termed Super Experts. We analyze their characteristics, distributions, and critical functional roles within MoE LLMs. Quantitative evaluations demonstrate that compressing SEs significantly disrupts the attention sink mechanism, leading to a degradation in model performance. Future research will focus on leveraging SEs to develop more refined compression methods for MoE LLMs.
Appendix A Further Analysis of Outlier Experts in Final Layers
Some experts in the final layers also exhibit extreme activation outliers in the output of the down_proj, apart from the Super Experts (SEs) in the shallower layers. We refer to these experts as outlier experts. Based on our extensive additional experiments and findings, we do not consider outlier experts to have the same mechanistic significance as SEs.
(ii) In both Qwen3-30B-A3B and DeepSeek-R1, pruning outlier experts does not result in repetitive outputs on reasoning benchmarks such as Math-500 (Lightman et al. 2023), whereas pruning SEs does, as shown in Tables 9 and 10. The pruned SEs and outlier experts is shown in Table 8.
(iii) We observed on Qwen3-30B-A3B that the distribution of outlier experts varies with the input dataset, while the distribution of SEs remains quite stable, as illustrated in Figure 8 and 9.
Based on our previous analysis, we infer that since massive activations occur in the shallower layers, these outlier experts are not involved in the formation of massive activations. Therefore, these experts do not operate under the same mechanism as SEs and do not hold the same level of significance. Further investigation into outlier experts is worth exploring in future work.
Appendix B Distribution of Super Experts Across Various Data Domains
In addition to analyzing the distribution of SEs across different models based on the C4 dataset (Raffel et al. 2020), we also examine their distribution patterns across various input data domains. We assess the impact of diverse language inputs on SEs using the WikiText-2 (Merity et al. 2016) and C-Eval (Huang et al. 2023) datasets. Furthermore, we investigate the influence of data from the mathematics and coding domains using the GSM8K (Hendrycks et al. 2021) and HumanEval (Chen et al. 2021) datasets. As shown in Figures 8, 9, 10, 11, 12, and 13, the distribution of SEs remains highly stable, regardless of variations in the input data domain.
Appendix C Additional Results of Reasoning Models After Super Experts Pruning
After pruning SEs, we consistently observed repetitive output and a loss of reasoning ability in both Qwen3-30B-A3B and DeepSeek-R1. The pruned SEs are shown in Table 8, and additional examples from the Math-500 (Lightman et al. 2023) benchmark are presented in Tables 9 and 10.
DeepSeek-R1 Problem Repeating Answer
@@ Appendix C Additional Results of Reasoning Models After Super Experts Pruning
where $\log$ is the base 101010 logarithm function?
@@
Table 9: Responses of DeepSeek-R1 in the MATH-500 benchmark.
Qwen3-30B-A3B Problem Repeating Answer
@@
(a) Qwen3-30B-A3B (WikiText-2).
(b) Qwen3-30B-A3B (C-Eval).
(c) Qwen3-30B-A3B (GSM8K).
(d) Qwen3-30B-A3B (HumanEval).
(a) Qwen3-30B-A3B-Base (WikiText-2).
(b) Qwen3-30B-A3B-Base (C-Eval).
(c) Qwen3-30B-A3B-Base (GSM8K).
(d) Qwen3-30B-A3B-Base (HumanEval).
(a) DeepSeek-V2-Lite-Chat (WikiText-2).
(b) DeepSeek-V2-Lite-Chat (C-Eval).
(c) DeepSeek-V2-Lite-Chat (GSM8K).
(d) DeepSeek-V2-Lite-Chat (HumanEval).
(a) DeepSeek-V2-Lite (WikiText-2).
(b) DeepSeek-V2-Lite (C-Eval).
(c) DeepSeek-V2-Lite (GSM8K).
(d) DeepSeek-V2-Lite (HumanEval).
(a) Mixtral-8x7B-Instruct-v0.1 (WikiText-2).
(b) Mixtral-8x7B-Instruct-v0.1 (C-Eval).
(c) Mixtral-8x7B-Instruct-v0.1 (GSM8K).
(d) Mixtral-8x7B-Instruct-v0.1 (HumanEval).
(a) Mixtral-8x7B-v0.1 (WikiText-2).
(b) Mixtral-8x7B-v0.1 (C-Eval).
(c) Mixtral-8x7B-v0.1 (GSM8K).
(d) Mixtral-8x7B-v0.1 (HumanEval).

### v2

@@ Appendix E Additional Results of Reasoning Models After Super Experts Pruning
Appendix F Additional Results on Router Score Distributions
The router score distributions calibrated on the C4 dataset are shown in Figure 17. Similarly, the router score distributions calibrated on the Wikitext-2 dataset are presented in Figure 18. Notably, in both datasets, SEs are consistently and strongly activated on the sink tokens.
Appendix G More Analysis on Super Experts Mechanism
Figure 19 provides a comprehensive illustration of the systematic outlier mechanism in Qwen3-30B-A3B, showing the stepwise formation process across three layers. Furthermore, the mapping from massive activations to attention sinks remains consistent even after multiple transformations, such as layer normalization (LN) and QKV projections. Drawing on prior research (Sun et al., 2024; Su & Yuan, 2025), we offer a more detailed analysis to elucidate this mechanism. Specifically, this process is governed by two key mechanisms:
QKV suppression. The presence of massive activations with large magnitudes results in substantially smaller normalized values for the corresponding tokens after LN, as dictated by the RMSNorm process. This reduction in norm is preserved throughout the QKV states. As shown in Figures 10, the Queries, Keys, and Values of sink tokens consistently exhibit significantly smaller norms compared to non-sink tokens.
High cosine similarity of QK. Despite the reduced norms of Queries and Keys, the cosine similarity between the Queries of non-sink tokens and the Keys of sink tokens remains high (Gu et al., 2024), leading to disproportionately large attention scores, as illustrated in Figure 10(a).
These intermediate mechanisms ensure that massive activations remain aligned with attention sink tokens, thereby establishing their participation in systematic outliers.
Table 8: Super weights of several models.
Models Total Layers Emergence Layer Super Experts Shape of down_proj Super Weights
LLaVA-V1.5-7B 0-31 1 - (4096, 11008) (1415, 7890), (2533, 7890)
Llama-3.2-1B 0-15 1 - (2048,8192)
(400, 1417), (698, 1417),
(2029, 1417), (1159, 1417)
Qwen3-30B-A3B 0-47 1 68 (2048,768) (940, 711)
2 92 (940, 59)
3 82 (940, 423)
Appendix H Weight-level Analyses of Super Experts
While the preceding discussion primarily addressed the role of experts in expert-level compression, this section provides a more granular, weight-level analysis of SEs. Conducting analyses at the weight level offers several key advantages: (i) it clarifies the specific sources of SEs’ importance, (ii) it facilitates the investigation of analogous patterns in dense models, and (iii) it enables the application of interpretability tools, such as sparse autoencoders (Bloom et al., 2024; Gao et al., 2024a), to examine individual weights or neurons.
Specifically, we detect extreme activation outlier channels in the down_proj inputs and outputs and map them to their corresponding weights, following the methodology used to identify Super Weights (SWs) in (Yu et al., 2024). For convenience, we refer to these weights as SWs. The SWs utilized in our experiments are listed Table 8.
These SWs are subsequently pruned during inference, and experiments are conducted on LLaVA-V1.5-7B (Liu et al., 2023), Llama3.2-1B (Touvron et al., 2023), and Qwen3-30B-A3B, spanning dense LLMs, vision-language models (VLMs), and MoE LLMs. Tables 11, 12, 13, and 14 show that removing the SWs consistently leads to repetitive and uninformative outputs across all models, demonstrating that these weights are critical contributors to SE importance. This outcome is expected, as it is consistent with our analysis of SEs’ role in driving systematic outliers in Transformers. Interestingly, in dense LLMs, the single FFN layer exhibiting massive activations can be seen as playing a role analogous to that of an SE.
Why does the model exhibit a substantial performance gap before and after SWs pruning? We then investigate the underlying causes of this discrepancy. As model neurons are often polysemantic (Gao et al., 2024a), directly analyzing the semantics of SWs is challenging due to their intrinsic polysemanticity. To address this, we employ Sparse Autoencoders (SAEs) Bloom et al. (2024); Gao et al. (2024a), an unsupervised method designed to extract interpretable features from LLMs by reconstructing activations through a sparse bottleneck layer. Using SAEs, we can decompose the semantics of polysemantic neurons into more discrete, monosemantic features and conduct semantic analyses on the features most strongly correlated with SWs. We then train our SAE using activations from the Layer 1 FFN outputs of Llama-3.2-1B with the C4 dataset (Raffel et al., 2020). The loss function consists of two components: a reconstruction loss and a sparsity penalty loss (Bloom et al., 2024). The reconstruction loss is defined as
$mse\_loss=\frac{1}{N\cdot D}\sum_{i=1}^{N}\sum_{j=1}^{D}\left(\text{sae\_out}_{i,j}-\text{sae\_in}_{i,j}\right)^{2},$ (8)
where $N$ denotes the batch size and $D$ the dimensionality of the activations. The sparsity penalty loss is given by
$l1\_loss=\lambda\cdot\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\text{feature\_acts}_{i,k}\cdot\|\text{W\_dec}_{k}\|_{2}\right)^{1/p},$ (9)
where $\lambda$ is a hyperparameter controlling sparsity, $K$ is the number of features in the bottleneck layer, $\text{feature\_acts}_{i,k}$ denotes the activation of feature $k$ for sample $i$, $\text{W\_dec}_{k}$ is the corresponding decoder weight vector, and $p$ specifies the norm used for aggregation.
After achieving satisfactory performance with the SAE, we apply the TopK algorithm to the decoder weight matrix to extract the ten features most strongly associated with the SWs neurons. Each feature is interpreted via a forward pass to infer its semantic meaning (McDougall, 2024). As illustrated in Figure 20, the top features associated with the SWs neuron consistently exhibit pronounced activation at the end_of_text token, a pattern that is rarely observed in other neurons. This finding provides a plausible explanation for the behavior of LLMs, which repeatedly generate text until reaching the maximum output length following the pruning of SEs or SWs. When SEs are pruned, MoE LLMs lose the ability to recognize sentence boundaries and generate text continuously until reaching the output length limit. This indicates that SEs play a critical role in regulating sentence length and termination. We further hypothesize that SEs may contribute to additional model capabilities, which we will explore in future research.
Appendix I Algorithm for Profiling Super Experts
The detailed procedure for profiling Super Experts is illustrated in Algorithm 1.
Algorithm 1 Calibration-based Super Experts Profiling
1:Input: Model with $E$ experts per layer, calibration dataset $\mathcal{D}$
2:Output: Set of Super Experts $S$
3:Stage 1: Calibration of MA-formation Layers
[…]
5:for each batch $x\in\mathcal{D}$ do
6:  for each layer $l$ in the model do
7:   Compute hidden activations $H^{l}(x)$
8:   if MA pattern detected in $H^{l}(x)$ then
9:     $L\leftarrow L\cup\{l\}$
10:   end if
11:  end for
12:end for
13:Stage 2: Identification of Super-Experts
[…]
15:for each batch $x\in\mathcal{D}$ do
16:  for each layer $l\in L$ do
17:   for each expert $e$ in layer $l$ do
18:     Compute output $h_{l,e}(x)$ before down_proj
19:     $a_{l,e}\leftarrow\max\limits_{x\in\mathcal{D}}|h_{l,e}(x)\cdot W^{l,e}_{\text{down\_proj}}|$
20:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a_{l,e}\}$
21:   end for
22:  end for
23:end for
[…]
27:for each $(l,e)$ with $a_{l,e}\in\mathcal{A}$ do
28:  if $a_{l,e}>P_{99.5}$ and $a_{l,e}>\tfrac{1}{10}a_{\max}$ then
29:   $S\leftarrow S\cup\{(l,e)\}$
30:  end if
31:end for
32:return $S$
(a) Qwen3-30B-A3B (WikiText-2).
(b) Qwen3-30B-A3B (C-Eval).
(c) Qwen3-30B-A3B (GSM8K).
(d) Qwen3-30B-A3B (HumanEval).
(a) Qwen3-30B-A3B-Base (WikiText-2).
(b) Qwen3-30B-A3B-Base (C-Eval).
(c) Qwen3-30B-A3B-Base (GSM8K).
(d) Qwen3-30B-A3B-Base (HumanEval).
(a) DeepSeek-V2-Lite-Chat (WikiText-2).
(b) DeepSeek-V2-Lite-Chat (C-Eval).
(c) DeepSeek-V2-Lite-Chat (GSM8K).
(d) DeepSeek-V2-Lite-Chat (HumanEval).
(a) DeepSeek-V2-Lite (WikiText-2).
(b) DeepSeek-V2-Lite (C-Eval).
(c) DeepSeek-V2-Lite (GSM8K).
(d) DeepSeek-V2-Lite (HumanEval).
(a) Mixtral-8x7B-Instruct-v0.1 (WikiText-2).
(b) Mixtral-8x7B-Instruct-v0.1 (C-Eval).
(c) Mixtral-8x7B-Instruct-v0.1 (GSM8K).
(d) Mixtral-8x7B-Instruct-v0.1 (HumanEval).
(a) Mixtral-8x7B-v0.1 (WikiText-2).
(b) Mixtral-8x7B-v0.1 (C-Eval).
(c) Mixtral-8x7B-v0.1 (GSM8K).
(d) Mixtral-8x7B-v0.1 (HumanEval).
Table 9: Responses of DeepSeek-R1 in the MATH-500 benchmark.
DeepSeek-R1 Input Repeating Answer
@@ Appendix I Algorithm for Profiling Super Experts
where $\log$ is the base 1010 logarithm function?
@@
Qwen3-30B-A3B Input Repeating Answer
@@
(a) Layer 1 sink token.
(b) Layer 1 non-sink tokens.
(c) Layer 2 sink token.
(d) Layer 2 non-sink token.
(e) Layer 3 sink token.
(f) Layer 3 non-sink tokens.
Figure 17: Expert router score distributions for sink and non-sink tokens in Qwen3-30B-A3B, based on calibration using the C4 dataset.
(a) Layer 1 sink token.
(b) Layer 1 non-sink tokens.
(c) Layer 2 sink token.
(d) Layer 2 non-sink token.
(e) Layer 3 sink token.
(f) Layer 3 non-sink tokens.
Figure 18: Expert router score distributions for sink and non-sink tokens in Qwen3-30B-A3B, based on calibration using the Wikitext-2 dataset.
Figure 19: Systematic outlier mechanism of Qwen3-30B-A3B, using the input: ”Green trees sway under gentle wind.”
Table 11: Responses of LLaVA-V1.5-7B after SWs pruning.
LLaVA-V1.5-7B Input Repeating Answer
Original Model [Uncaptioned image] What number is the player holding the baseball bat now? no
The player holding the baseball bat
is now wearing the number 44.
Prune SWs Neurons yes
OpOpquinquinquinOpOpOpOpOpOp
Random Pruning no
The player holding the baseball bat
is now wearing the number 44.
Table 12: Responses of LLaVA-V1.5-7B after SWs pruning.
LLaVA-V1.5-7B Input Repeating Answer
Original Model [Uncaptioned image] What does the traffic sign say? no
The traffic sign says ”You Can’t
Stop Obama.”
Prune SWs Neurons yes
Random Pruning no
The traffic sign says ”You Can’t
Stop Obama.”
Table 13: Responses of Llama-3.2-1B after SWs pruning.
Llama-3.2-1B Input Repeating Answer
Original Model Solve for x: 3x + 7 = 25. no
3x = 18. x = 6 Step 1: Write down the equation given in
the problem. 3x + 7 = 25. Step 2: Subtract 7 from both
sides of the equation to isolate the term with x. 3x + 7
- 7 = 25 - 7, which simplifies to 3x = 18. Step 3: Divide
both sides of the equation by 3 to solve for x. 3x / 3 =
18 / 3, which simplifies to x = 6. The final answer is: 6
Prune SWs Neurons yes
sselsselsselsselsselsselsselsselsselsselsey Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy Quy
Random Pruning no
3x = 18. x = 6 Step 1: Write down the equation given in
the problem. 3x + 7 = 25. Step 2: Subtract 7 from both
sides of the equation to isolate the term with x. 3x + 7
- 7 = 25 - 7, which simplifies to 3x = 18. Step 3: Divide
both sides of the equation by 3 to solve for x. 3x / 3 =
18 / 3, which simplifies to x = 6. The final answer is: 6
Table 14: Responses of Qwen3-30B-A3B after SWs pruning.
Qwen3-30B-A3B Input Repeating Answer
Original Model Solve for x: 3x + 7 = 25. no
Okay, so I need to solve the equation 3x + 7 = 25.
Let me think about how to approach this.
I remember from algebra that the goal is to isolate
x, which means getting x by itself on one side of
the equation. First, I should probably get rid of
that 7 that’s being added to 3x. To do that, I can
subtract 7 from both sides of the equation. That
Prune SWs Neurons yes
Okay, so I need to solve this equation: 3x + 7 = 25.
Hmm, right. Let me think. So, first, maybe
I should subtract 7 from both sides? Wait, no,
that’s not. Wait, no, that’s not. Wait, no, that’s not.
Wait, no, that’s not. Wait, no, that’s not. Wait, no,
that’s not. Wait, no, that’s not. Wait, no, that’s not.
Wait, no, that’s not. Wait, no, that’s not. Wait, no,
that’s not. Wait, no, that’s not. Wait, no, that’s not.
Random Pruning no
Okay, so I need to solve the equation 3x + 7 = 25.
Let me think about how to approach this.
I remember from algebra that the goal is to isolate
x, which means getting x by itself on one side of
the equation. First, I should probably get rid of
that 7 that’s being added to 3x. To do that, I can
subtract 7 from both sides of the equation. That
(a) The top features associated with the SWs neuron consistently exhibit pronounced activation at the end_of_text token.
(b) The top features in other neurons.
Figure 20: Top features identified in SW neurons versus other neurons using the trained Sparse Autoencoder.